A container has ‘a’ defective and 3 non defective items. if a sample of 2 items

are drawn one after another without replacement what is the probability that, the

sample will at most one defective


1
Expert's answer
2021-08-10T06:44:58-0400

Let "x=" the number of defective items among 2 drawn items.

"n=a+3"


"P(x\leq1)=P(x=0)+P(x=1)"


"=\dfrac{\dbinom{a}{0}\dbinom{3}{2}}{\dbinom{a+3}{2}}+\dfrac{\dbinom{a}{1}\dbinom{3}{1}}{\dbinom{a+3}{2}}"

"=3\dfrac{2!(a+1)!}{(a+3)!}+3a\dfrac{2!(a+1)!}{(a+3)!}"

"=\dfrac{6(a+1)}{(a+2)(a+3)}"

The probability that, the sample will at most one defective is "\dfrac{6(a+1)}{(a+2)(a+3)} ."